OpenStudy (anonymous):

For a rectangle with a perimeter 44 to have the largest area, what dimensions should it have?

OpenStudy (ash2326):

let the length be l and breadth be b perimeter=2(l+b)=44 so l+b=22 now Area A=l*b in terms of l b=22-l so A=l*(22-l) now differentiate A w.r.t l dA/dl=22-2l for finding the value of l for which area is max, equate dA/dl to 0 22-2l=0 l=11 this is the value of l for which area is max b=22-l=22-11=11 so l =11 and b=11 area=121 unit^2